Circular No. 8436

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 TU_{12} (SIDING SPRING) An apparently asteroidal object discovered on Oct. 10 by the Siding

Spring Survey (announced as 2004 TU_{12} on MPEC 2004-T55; discovery observation below) has been found to show a short eastward tail. CCD images taken over 2.5 hr with the 0.36-m “SoTIE” reflector at Las Campanas by F. Mallia, G. Masi, and R. Wilcox on Nov. 12.0 UT show a tail ∼ 2′ long in p.a. 70^{o} (the “head” appearing like stars in the field (FWHM ∼ 3′′). Eighty 30-s CCD images taken on Nov. 12.8 by J. Lacruz (La Cañada, Spain) show a sharp tail 4′ long in p.a. 69^{o} (and total mag 14.3).

2004 UT  α_{2000}  δ_{2000}  Mag.
Oct. 10.55030  22^{h}52^{m}33.^{s}67  −51^{o}01′39′′.8  14.1

The following orbital elements from MPEC 2004-V56 are from identified observations spanning 1990–2004 (H_{10} = 15.0).

Epoch = 2005 Jan. 30.0 TT

T = 2004 Nov. 10.2817 TT  ω = 356.^{o}3614
e = 0.597279  Ω = 31.2525 }2000.0
q = 1.227632 AU  i = 27.8448

a = 3.048339 AU n^{o} = 0.1851864 P = 5.32 years

2004TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  Mag.
Nov. 11 22^{h}50.^{m}63 −11^{o}25.^{′}4 0.456 1.228 110.^{o}7 49.^{o}0 14.2
21 23 01.38 + 1 29.3 0.495 1.234 108.0 49.6 14.4
Dec. 1 23 15.80 +12 33.8 0.555 1.252 105.2 49.4 14.7
11 23 33.48 +21 38.5 0.631 1.281 102.7 48.6 15.1
21 23 54.35 +28 58.6 0.718 1.319 100.5 47.2 15.5
31 0 18.32 +34 54.0 0.814 1.366  98.5 45.4 15.9

SUPERNOVA 2004fu IN NGC 6949 M. Modjaz, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 350–740 nm) of SN 2004fu (cf. IAUC 8428), obtained by M. Westover on Nov. 10.17 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova, several days before maximum light. Adopting the NED recession velocity of 2761 km/s for the host galaxy, the ejecta expansion velocities are 14000 km/s for Si II (rest 635.5 nm) and 22700 km/s for Ca II (rest 395.1 nm).

2004 November 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green